How many molecules of Ag2So4 will be made if 2.7 grams of AgNO3 are reacted?
AgNO3 + H2SO4 --> Ag2SO4 + HNO3
1
Expert's answer
2020-12-11T02:51:21-0500

Solution:

The balanced chemical equation:

2AgNO3 + H2SO4 --> Ag2SO4 + 2HNO3

According to the equation: n(AgNO3) / 2 = n(Ag2SO4)


Moles of AgNO3 = n(AgNO3) = Mass of AgNO3 / Molar mass of AgNO3

The molar mass of AgNO3 is 169.87 g/mol.

Hence,

n(AgNO3) = 2.7 g / 169.87 g mol-1 = 0.01589 moles


n(Ag2SO4) = n(AgNO3) / 2 = 0.01589 moles / 2 = 0.007945 moles


One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

Number of Ag2SO4 molecules = (0.007945 mol) × (6.022×1023 molecules / 1 mol) = 4.78×1021 molecule

Number of Ag2SO4 molecules = 4.78×1021 molecules


Answer: 4.78×1021 molecules of Ag2SO4 will be made.


The final calculation formula considering all steps:

2.7 g AgNO3 × (1 mol AgNO3 / 169.87 g AgNO3) × (1 mol Ag2SO4 / 2 mol AgNO3) × (6.022×1023 molecules Ag2SO4 / 1 mol Ag2SO4) = 4.78×1021 molecules Ag2SO4